The sum of 8 and a certain number is equal to the product of the number and 3. find the number
Found 2 solutions by mathstutor494, Edwin McCravy:
Answer by mathstutor494(120)

Let a certain number be "x"
sum of 8 and a certain number = x+8
product of the number and 3 = 3x
Given that above 2 statements are equal, so
x+8= 3x
3x-x = 8
2x = 8
so x=4........................Ans

Answer by Edwin McCravy(20060)

Here it is step by step:

The sum of 8 and a certain number is equal to the product of the number and 3.
Replace the words "a certain number" and "the number" by N.  Then
the sentence becomes:

The sum of 8 and N is equal to the product of N and 3.
Replace the words "The sum of 8 and N" by 8+N.  Then
the sentence becomes:

8+N is equal to the product of N and 3.
Replace the words " the product of N and 3" by N×3.  Then
the sentence becomes:

8+N is equal to N×3.
Replace the words "is equal to" by the symbol =, an equal sign.
Then the sentence becomes:

8+N = N×3
To avoid the × sign for multiplication we use the commutative principle
and replace N×3 by 3N.  Then the sentence becomes:

8+N = 3N
Then we subtract N from both sides and get

8 = 2N
Then we divide both sides by 2 and get:

4 = N
So the number is 4.


Edwin